OpenStudy (anonymous):

The speed of sound in water is 1498 m/s. A sonar signal is sent from a ship at a point just below the water surface and 1.80 s later the reflected signal is detected. How deep is the ocean beneath the ship? uggghh

OpenStudy (anonymous):

s=vxt/2=1348.2

OpenStudy (jamesj):

By definition Speed = (Distance) / (Time) thus Distance = (Speed)(Time). In this case Speed = 1,498 m/s Time = 1.80 s Therefore, in this case, the sound has travelled a distance of d = (speed)(time) = (1498 m/s)(1.80 s) = 2,696.4 m Now, if the sound has travelled that far, how far away is the the bottom of the ocean?